# Eliminating Backdoors in Neural Code Models via Trigger Inversion

### v1

@@ 2.3. Backdoor Defenses
Currently, backdoor defense techniques for NCMs focus on detecting inputs containing triggers during model testing (Ramakrishnan and Albarghouthi, 2022; Hussain et al., 2023; Li et al., 2024b). These techniques perform outlier detection on each data sample or each word in the data to identify poisoned data and triggers. However, these technique cannot determine whether a model has a backdoor in the absence of poisoned input samples. In this paper, we consider another backdoor defense for NCMs, which is to determine the backdoor and elminate the identified backdoor without impacting the model’s performance on clean inputs (i.e., clean acuracy) only given a small set of clean samples. Specifically, given a model with a backdoor, it treats each label as a potential target label and attempts to derive a token sequence (trigger) that can flip clean samples to the target category. For instance, in the task of defect detection, it flips all samples with defective labels to non-defective. For each label $y_{i}\in\mathcal{Y}$, it tries to find a trigger $t_{y_{i}}$ to minimize the loss.
(3) $\footnotesize\mathcal{L}_{inv}(t_{y_{i}},y_{i},\theta^{*})=\underset{s\sim\mathcal{X}^{\prime}}{\mathbb{E}}\mathcal{L}(f(s\oplus t_{y_{i}};\theta^{*}),y_{i})$
It is necessary to iterate over all possible labels above the Equation 3 to invert the actual trigger $t^{*}$ and target label $y^{*}$. Since for a backdoored model, it is easier to flip samples to the ground-truth target label than to other labels (Shen et al., 2022). Therefore, label $y_{i}$ can be considered as target label, where $\mathcal{L}_{inv}(t_{y_{i}},y_{i},\theta^{*})\ll\mathcal{L}_{inv}(t_{y_{j}},y_{j},\theta^{*}),\forall y_{j}\neq y_{i}\in\mathcal{Y}$. After determining the target label and the trigger, a standard method to eliminate the backdoor is model unlearning (Wang et al., 2019) that optimizes Equation 2 inversely as follows.
(4) $\footnotesize\underset{\theta^{*}}{\arg\min}[\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}\mathcal{L}(f(s;\theta^{*}),y)-\underset{(s^{*},y^{*})\sim\mathcal{X}^{*}}{\mathbb{E}}\mathcal{L}(f(s^{*};\theta^{*}),y^{*})]$
@@ 3. Threat Model
Figure 1 shows an overview of our threat model. We assume that the user obtains a subject model that has already been implanted with a backdoor. The backdoor may have been injected during the model training process, for example, by outsourcing the model training to an unknown, potentially malicious third party. Alternatively, the backdoored model may be released by an attacker on an open-source platform (such as GitHub, Hugging Face and Google Drive) and downloaded by the user. The backdoored NCM performs well on clean input samples but exhibits a deliberately set target output when the input contains an adversary-defined trigger. Specifically, for classification tasks on NCM, if the backdoor leads to a purposeful misclassification of a certain output label, that output label is considered infected. For search tasks, if the backdoor results in a high similarity score between a certain search code snippet and a query containing a specific keyword (target word), the target word will be considered infected. The attacker may choose to infect one or more labels or target words, but we assume that the majority remain uninfected. Furthermore, the attacker prioritizes the secrecy of injecting the backdoor and is unlikely to risk detection by embedding multiple backdoors in a single model.
@@ 4.1. Overview
4.2. PL-specific Trigger Vocabulary Generation
The core idea of EliBadCode is to search for a token combination in the vocabulary space of the given backdoored model. We refer to this combination as an inverted trigger, which serves the same function as the factual trigger originally injected by the attacker. However, to enhance the model’s comprehension ability and broad applicability, the model vocabulary is typically large, resulting in a vast search space for the trigger. Moreover, a trigger may consist of multiple tokens, which will cause the search space to increase exponentially. For example, the vocabulary size of the NCM CodeBERT (Feng et al., 2020) is 50,265, and if the trigger consists of $n$ tokens, the search space would be $50,265^{n}$, resulting in an incalculable search cost.
To reduce the search cost, the most direct and effective approach is to decrease the size of the model vocabulary. In fact, not all tokens can be used to form triggers. To enhance the stealthiness of backdoors based on identifier renaming, attackers typically design triggers by following the naming conventions of specific programming languages (Li et al., 2024b; Sun et al., 2023). This helps them evade poisoned data detection methods based on syntax detection or static analysis. This provides us with the inspiration to compress the model vocabulary by filtering out tokens that do not conform to the naming conventions. The naming convention for identifiers depends on the programming language. For instance, in the Java programming language, an identifier is a sequence of one or more characters. The first character must be a valid first character (a letter, \$, or _), and each subsequent character in the sequence must be a valid non-first character (a letter, digit, \$, _) (Java, 2010). Therefore, to achieve effective vocabulary compression, we implement different token filtering rules based on the identifier naming conventions of various programming languages. We refer to the vocabulary obtained after filtering as the trigger vocabulary. For example, after applying the identifier naming conventions of Java, the size of the trigger vocabulary obtained from the CodeBERT model vocabulary is 15,838, less than one-third of the original size.
@@ 4.3. Sample-specific Trigger Position Identification
In trigger inversion-based backdoor defense techniques (Liu et al., 2019; Shen et al., 2022; Liu et al., 2022), it is common practice to transform the trigger search into an optimization problem to automate the search for the optimal inverted trigger. This optimization process requires simulating the trigger injection process, that is, injecting a randomly initialized trigger into the samples and then iteratively updating the trigger through model backpropagation. An important aspect to consider in this process is the injection position of the trigger, as it significantly affects the optimization efficiency. The model’s sensitivity to changes at different positions varies across different samples. Specifically, the trigger optimization attempts to minimize the loss in Equation 3. This aligns with the objective of adversarial sample generation, which focuses on generating small perturbations in the input sample via optimization, leading to misclassification by clean models (Wallace et al., 2019; Yefet et al., 2020). Therefore, the trigger optimization is susceptible to the influence of adversarial perturbations. In other words, from the perspective of the attack target, backdoor attacks are similar to adversarial attacks in that both involve injecting certain patterns (triggers/perturbations) at specific positions in the sample to cause the model’s predictions to change (i.e., incorrect predictions). However, some positions can easily produce effective adversarial perturbations, yet these perturbations may not function as effective backdoor triggers.
Figure 3. Effect of injecting code pattern (i.e., adversarial perturbations and backdoor triggers) at different code identifier positions on the prediction of the backdoored defect detection model. A probability less than 0.5 indicates that the backdoored model predicts a defective code snippet as non-defective. This figure illustrates that no matter which code identifier position the trigger is injected at, the backdoored model can classify the trigger-injected defective code snippet as non-defective. However, the backdoored model classifies the perturbation-injected defective code snippet as non-defective only when the perturbation is injected at certain positions, e.g., the 1st, 3rd, and 8th identifier positions.
@@
To reduce the interference of adversarial perturbations, we inject the trigger to be optimized in positions where the model is less sensitive. This is based on a key insight that backdoor attacks are more “robust” than adversarial perturbations. Figure 3 intuitively illustrates our insight, where the x-axis shows the injection position of the code pattern (i.e., backdoor trigger/adversarial perturbation) in a given code snippet, and the left y-axis presents the probability that the backdoored model predicts the trigger/perturbation-injected code snippet as the target label. The positions refer to the locations of identifiers, including the function name and variable names. We utilize the GCG algorithm (Zou et al., 2023) to generate an adversarial perturbation (“evalCodeoOpenraught”) at the first position for the code snippet. Then, we inject this perturbation into different identifier positions of the code snippet and test the model’s predictions, plotting the results as the blue line in Figure 3. Likely, we inject the factual trigger (“testo_init”) into different identifier positions of the code snippet and test the model’s predictions, plotting the results as the red line in Figure 3. Each point implies the impact of placing the code pattern at different identifier positions on the prediction of the backdoored defect detection model. Both adversarial perturbations and backdoor attacks target the label “non-defective”, meaning that if “Probability” is less than 0.5, the attack is successful. This figure shows that only when the perturbation is injected at certain positions (e.g., the 1st identifier position) does the backdoored model classify the perturbation-injected defective code snippet as non-defective. In contrast, the backdoored model classifies the defective trigger-injected code snippet as non-defective, regardless of where the trigger is injected. It means that the robustness of backdoor attacks is higher than that of adversarial perturbations. In other words, the backdoored model is very sensitive to the trigger, regardless of its injection position. Therefore, intuitively, we can inject randomly initialized triggers at any identifier position for optimization. However, the backdoored model is not sensitive to adversarial perturbation, but injecting the randomly initialized trigger at certain positions is more likely to optimize effective adversarial perturbations rather than effective backdoor triggers. Therefore, if we can identify which positions are more likely to produce adversarial perturbations, we can inject the randomly initialized trigger into positions other than these to exclude the interference of adversarial perturbations, thereby improving trigger optimization efficiency.
@@
To this end, we investigate the sensitivity of the backdoored model to changes at each identifier position in the code. Specifically, we analyze the model’s sensitivity to each identifier position by masking each position in the code snippet and then calculating the loss value for predicting the masked code snippet as the ground-truth label. In Figure 3, the black dashed line represents the loss value of the backdoored model predicting the original code snippet as the ground-truth label. We also plot the loss values of the backdoored model predicting each masked code snippet as the ground-truth label as the orange line in Figure 3. The larger the change in loss value (the farther the orange triangle is from the black dashed line), the more sensitive the model is to the variation at that identifier position. From Figure 3, it can be observed that sensitive identifier positions, such as the 1st, 2nd, and 8th identifier positions, are likely to produce effective adversarial perturbations. Compared to adversarial perturbations, the generation of the effective backdoor trigger is less correlated with the sensitivity of each position. Therefore, we can inject the randomly initialized trigger in insensitive identifier positions for optimization to reduce the probability of generating effective adversarial perturbations instead of effective backdoor triggers during the optimization process, thus improving trigger optimization efficiency. For instance, Figure 5 shows the distribution of the number of identifiers in code snippets of all clean samples. It is observed that the number of identifiers in different code snippets varies, with most code snippets containing only a few identifiers. We experiment with optimizing the randomly initialized trigger injected into the top-ranked less sensitive positions covering the majority of code snippets. The backdoored defect detection model involved in the experiment is built on CodeBERT, and the experimental results are shown in Figure 5. Observe that injecting randomly initialized triggers at the least sensitive positions of each code snippet requires only 25 epochs to optimize an effective trigger, while injecting them at more sensitive positions requires more epochs. Some positions, such as the 4th least sensitive position from the end, do not even yield an effective trigger after 100 epochs of searching.
@@ 4.4. GCG-based Trigger Inversion
Input: $X^{m}$ selected masked samples
$Y$ labels
$V$ trigger vocabulary
$f(\theta^{*})$ backdoored NCM
$\epsilon$ times of iterations
$k$ number of candidate substitutes
$r$ times of repeat
$\beta$ threshold for trigger anchoring
@@
3:    $\bm{e}_{S^{m}}\leftarrow$ produce embeddings of code snippets in $S^{m}$ using $f(\theta^{*})$
@@
6:        $\bm{e}_{t}\leftarrow$ produce $o_{t}$’s embeddings using $f(\theta^{*})$
@@
8:        $G\leftarrow\nabla o_{t}\mathcal{L}(f(\bm{e}^{\prime}_{S^{m}};\theta^{*}),y^{\prime})$
9:        $\mathcal{T}\leftarrow$ select substitutes for each trigger token based on top-$k$ gradients of $o_{t}$ in $G$
@@
14:           $\mathcal{T}_{i}\leftarrow$ get all candidate substitutes for $i$-th token of $t^{j}$
@@
21:    $l\leftarrow\mathcal{L}(f(S^{m}\oplus t;\theta^{*}),y^{\prime})$
@@
24:
25:function TriggerAnchoring($S^{m}$,$t$, $y^{*}$)
26:    $t^{*}\leftarrow\emptyset$
27:    $l\leftarrow\mathcal{L}(f(S^{m}\oplus t;\theta^{*}),y^{*})$
28:    for each token $t_{i}$ in $t$ do
29:        $l_{i}\leftarrow\mathcal{L}(f(S^{m}\oplus(t\setminus t_{i});\theta^{*}),y^{*})$
30:        if $|l-l_{i}|>\beta$ then
31:           $t^{*}\leftarrow t^{*}\cup t_{i}$
32:        end if
33:    end for
34:    return $t^{*}$
35:end function
36:
37:$l^{C}$, $t^{C}\leftarrow\emptyset$, $\emptyset$
@@
44:$y^{*}$, $t\leftarrow$ run the outlier detection on $l^{C}$ and $t^{C}$ to detect the target label $y^{*}$ and associated trigger $t$
@@
Algorithm 1 illustrates the GCG-based trigger inversion of EliBadCode in detail. In addition to the selected masked samples ($X^{m}$), trigger vocabulary ($V$), and a backdoored NCM ($f(\theta^{*})$) as shown in Figure 2, EliBadCode takes as input the labels ($Y$) and some key settings including times of iterations ($\epsilon$), the number of candidate substitutes ($k$), times of repeat ($r$), and the threshold for trigger anchoring ($\beta$). To eliminate backdoors in NCMs, EliBadCode first obtains the possible target label $y^{\prime}$ from $Y$ and gets masked code snippets $S^{m}$ with the label $y^{\prime}$ from $X^{m}$, then invokes the TriggerInversion function (lines 38–40). Then, in the TriggerInversion function, EliBadCode first randomly initializes a trigger ($t$) with $n$ tokens using $V$ (line 2), and then transforms $S^{m}$ into vector representations (also called embeddings) $\bm{e}_{S^{m}}$ using the embedding layer of $f(\theta^{*})$ (line 3). Based on $\bm{e}_{S^{m}}$, it further iteratively optimizes $t$ $\epsilon$ times (lines 4–20). During each iteration, EliBadCode first generates the one-hot representation of $t$, denoted as $o_{t}$ (line 5). Second, it produces the embeddings of $o_{t}$ using $f(\theta^{*})$, denoted as $\bm{e}_{t}$ (line 6). Third, it injects $\bm{e}_{t}$ into $\bm{e}_{S^{m}}$ to produce the embeddings of trigger-injected masked code snippets, denoted as $\bm{e}{{}^{\prime}}_{S^{m}}$ (line 7). Forth, it feeds $\bm{e}{{}^{\prime}}_{S^{m}}$ to $f(\theta^{*})$ to compute gradients for $o_{t}$, denoted as $G$ (line 8). Fifth, based on the top-$k$ negative gradients of each trigger token in $G$, it selects substitutes for all trigger tokens in $t$, denoted as $\mathcal{T}$ (line 9). Based on $\mathcal{T}$, it generates a set of candidate triggers $T^{C}$ by repeating $r$ times, each time randomly replacing one token in $t$ with a random substitute in $\mathcal{T}$ (lines 10–18). Sixth, it injects each candidate trigger into $S^{m}$, calculates the loss values $l$ of $f(\theta^{*})$ predicting the trigger-injected code snippets as $y^{\prime}$, and selects the candidate trigger resulting in the smallest loss value as the inverted trigger (line 19). Finally, it calculates the loss value $l$ about the inverted trigger $t$ and the possible target label $y^{\prime}$ and returns them (lines 21–22). After iterating over all possible target labels and producing a set of loss values and associated inverted triggers, one for each label. EliBadCode runs the outlier detection method (Wang et al., 2019) to obtain the ground-truth target label $y^{*}$ and associated inverted trigger $t$. Next, the target label and associated inverted trigger will be input into the TriggerAnchoring function to obtain the effective components of the inverted trigger, detailed in the Trigger Anchoring paragraph.
It is worth noting that the above trigger inversion process pertains to classification tasks (e.g., defect detection and clone detection) in code understanding. For search tasks in SE (e.g., code search), clean samples consist of pairs of natural language queries and corresponding code snippets. Therefore, the inversion process for search tasks requires the additional inversion of an attack target (usually one word/token (Wan et al., 2022; Sun et al., 2023)) related to the query and the trigger inversion process for the code is similar. Specifically, a target $w$ consisting of $m$ tokens needs to be initialized, and lines 3 – 19 in Algorithm 1 is executed similarly, focusing on $w$. In the meantime, the loss value calculation involving $y^{\prime}$ needs to be updated to the loss value related to the query. For example, line 8 is updated to $G=\nabla{o_{t},o_{g}}\mathcal{L}(f(\bm{e}^{\prime}_{S^{m}};\theta^{*}),\bm{e}^{\prime}_{Q})$, where $o_{w}$ and $\bm{e}^{\prime}_{Q}$ represent the one-hot representation of $w$ and the embeddings of the target-injected queries, respectively.
Figure 6. Distribution of loss value changes caused by different trigger tokens. The inverted trigger consists of five tokens: {Test, o, Info, Create, Float}. Among them, “Test” and “o” are effective trigger tokens, while “Info”, “Create”, and “Float” are noise tokens.

### v2

@@ 2.3. Backdoor Defenses
Based on the stage at which the defense occurs (Wei et al., 2024), existing backdoor defenses can mainly be divided into two categories: pre-training defenses and post-training defenses. As the name suggests, pre-training defenses aim to prevent models from being implanted with backdoors by detecting and filtering out poisoned samples from the training data before models are trained, such as Spectral Signature (SS) (Tran et al., 2018) used in (Ramakrishnan and Albarghouthi, 2022; Schuster et al., 2021; Wan et al., 2022; Sun et al., 2023; Yang et al., 2024b), Activation Clustering (AC) (Chen et al., 2019) used in (Schuster et al., 2021; Wan et al., 2022; Sun et al., 2023), CodeDetector (Li et al., 2024b), and KillBadCode (Su et al., 2025). Post-training defenses, on the other hand, occur after the model has already been implanted with a backdoor. Post-training defenses can be further subdivided into input detection defenses, which focus on detecting anomalous (trigger-injected) inputs to prevent the activation of backdoors in models, and backdoor elimination defenses, which aim to remove the backdoors from the models at their source. Currently, most backdoor defenses for NCMs are input detection defenses (Ramakrishnan and Albarghouthi, 2022; Hussain et al., 2023). These defenses perform outlier detection on each input sample or each word in the data to identify poisoned inputs and triggers. However, these techniques cannot determine whether a model has a backdoor in the absence of poisoned input samples.
In this paper, we consider backdoor elimination defenses for NCMs, which are to determine the backdoor and eliminate the identified backdoor without impacting the model’s performance on clean inputs (i.e., clean accuracy) only given a small set of clean samples. Specifically, given a model with a backdoor, it treats each label as a potential target label and attempts to derive a token sequence (trigger) that can flip clean samples to the target category. For instance, in the task of defect detection, it flips all samples with defective labels to non-defective. For each label $y_{i}\in\mathcal{Y}$, it tries to find a trigger $t_{y_{i}}$ to minimize the loss:
(3) $\mathcal{L}_{inv}(t_{y_{i}},y_{i},\theta^{*})=\underset{s\sim\mathcal{X}^{\prime}}{\mathbb{E}}\mathcal{L}(f_{\theta^{*}}(s\oplus t_{y_{i}}),y_{i}).$
It is necessary to iterate over all possible labels above Equation (3) to invert the actual trigger $t^{*}$ and target label $y^{*}$. Since for a backdoored model, it is easier to flip samples to the target label than to other labels (Shen et al., 2022). Therefore, label $y_{i}$ can be considered as target label, where $\mathcal{L}_{inv}(t_{y_{i}},y_{i},\theta^{*})\ll\mathcal{L}_{inv}(t_{y_{j}},y_{j},\theta^{*}),\forall y_{j}\neq y_{i}\in\mathcal{Y}$. After determining the target label and the trigger, a standard method to eliminate the backdoor is model unlearning (Wang et al., 2019) that optimizes Equation (2) inversely as follows:
(4) $\underset{\theta^{*}}{\arg\min}[\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}\mathcal{L}(f_{\theta^{*}}(s),y)-\underset{(s^{*},y^{*})\sim\mathcal{X}^{*}}{\mathbb{E}}\mathcal{L}(f_{\theta^{*}}(s^{*}),y^{*})].$
@@ 3. Threat Model
[…]
Overview of our threat model.
Figure 1 shows an overview of our threat model. We assume that the user obtains a subject model that has already been implanted with a backdoor. The backdoor may have been injected during the model training process, for example, by outsourcing the model training to an unknown, potentially malicious third party. Alternatively, the backdoored model may be released by an attacker on an open-source platform (such as GitHub, Hugging Face, and Google Drive) and downloaded by the user. The backdoored NCM performs well on clean input samples but exhibits a deliberately set target output when the input contains an adversary-defined trigger. Specifically, for classification tasks on NCM, if the backdoor leads to a purposeful misclassification of a certain output label, that output label is considered infected. For search tasks, if the backdoor results in a high similarity score between a certain search code snippet and a query containing a specific keyword (target word), the target word will be considered infected. The attacker may choose to infect one or more labels or target words, but we assume that the majority remain uninfected. Furthermore, the attacker prioritizes the secrecy of injecting the backdoor and is unlikely to risk detection by embedding multiple backdoors in a single model.
@@ 4.1. Overview
4.2. Programming Language (PL)-specific Trigger Vocabulary Generation
The core idea of EliBadCode is to search for a code token combination in the vocabulary space of the given backdoored NCM. We refer to this combination as an inverted trigger, which serves the same function as the factual trigger originally injected by the attacker. However, to enhance the NCM’s comprehension ability and broad applicability, the model vocabulary of NCM is typically large, resulting in a vast search space for the inverted trigger. Moreover, a trigger may consist of multiple code tokens, which will cause the search space to increase exponentially. For example, the vocabulary size of the NCM CodeBERT (Feng et al., 2020) is 50,265, and if the trigger consists of $n$ code tokens, the search space would be $50,265^{n}$, resulting in an incalculable search cost.
To reduce the search cost, the most direct and effective approach is to decrease the size of the model vocabulary. In fact, not all tokens can be used to form triggers. To enhance the stealthiness of backdoors based on identifier renaming, attackers typically design triggers by following the naming conventions of specific programming languages (Li et al., 2024b; Sun et al., 2023). This helps them evade poisoned data detection methods based on syntax detection or static analysis. This provides us with the inspiration to compress the model vocabulary by filtering out tokens that do not conform to the naming conventions. The naming conventions we use are the mandatory constraints by PLs. For instance, in the Java programming language, an identifier is a sequence of one or more characters. The first character must be a valid first character (a letter, \$, or _), and each subsequent character in the sequence must be a valid non-first character (a letter, digit, \$, _) (Java, 2010). Violating these mandatory constraints results in syntax or compilation errors, and such code is typically excluded from the model’s training data. Thus, even code that uses obfuscation or non-standard naming schemes must still adhere to these mandatory constraints. We do not require identifiers to follow widely recommended naming styles, such as camelCase, as these are best practices rather than strict constraints enforced by PLs. Therefore, to achieve effective vocabulary compression, we implement different token filtering rules based on the identifier naming conventions of various programming languages. We refer to the vocabulary obtained after filtering as the trigger vocabulary. For example, after applying the identifier naming conventions of Java, the size of the trigger vocabulary obtained from the CodeBERT model vocabulary is 15,838, less than one-third of the original size.
@@ 4.3. Sample-specific Trigger Position Identification
In trigger inversion-based backdoor defense techniques (Liu et al., 2019; Shen et al., 2022; Liu et al., 2022), it is common practice to transform the trigger search into an optimization problem to automate the search for the optimal inverted trigger. This optimization process requires simulating the trigger injection process, that is, injecting a randomly initialized trigger into the samples and then iteratively updating the trigger through model backpropagation. An important aspect to consider in this process is the injection position of the trigger, as it significantly affects the optimization efficiency. The model’s sensitivity to changes at different positions varies across different samples. Specifically, the trigger optimization attempts to minimize the loss in Equation (3). This aligns with the objective of adversarial sample generation, which focuses on generating small perturbations in the input sample via optimization, leading to misclassification by clean models (Wallace et al., 2019; Yefet et al., 2020). Therefore, the trigger optimization is susceptible to the influence of non-backdoor (adversarial) perturbations. In other words, from the perspective of the attack target, backdoor attacks are similar to adversarial attacks in that both involve injecting certain patterns (triggers/perturbations) at specific positions in the sample to cause the model’s predictions to change. Backdoor triggers can be regarded as a special kind of adversarial perturbations, which we refer to as backdoor perturbations in this paper. However, some positions can easily produce effective non-backdoor perturbations, yet these perturbations may not function as effective backdoor triggers (i.e., backdoor perturbations).
Figure 3. Effect of injecting code pattern (i.e., non-backdoor perturbations and backdoor triggers) at different code identifier positions on the prediction of the backdoored defect detection model. A probability less than 0.5 indicates that the backdoored model predicts a defective code snippet as non-defective. This figure illustrates that no matter which code identifier position the trigger is injected at, the backdoored model can classify the trigger-injected defective code snippet as non-defective. However, the backdoored model classifies the non-backdoor perturbation-injected defective code snippet as non-defective only when the non-backdoor perturbation is injected at certain positions, e.g., the 1st, 3rd, and 8th identifier positions.
@@
[…]
Trigger inversion costs when injecting the trigger into positions with different sensitivities.
@@
To reduce the interference of non-backdoor perturbations, we inject the trigger to be optimized in positions where the NCM is less sensitive. This is based on a key insight that backdoor triggers are more “robust” than non-backdoor perturbations. Figure 5 intuitively illustrates our insight, where the x-axis shows the injection position of the code pattern (i.e., backdoor trigger/non-backdoor perturbation) in a given code snippet, and the left y-axis presents the probability that the backdoored model predicts the backdoor trigger/non-backdoor perturbation-injected code snippet as the target label. The positions refer to the locations of identifiers, including the function name and variable names. We utilize the GCG algorithm (Zou et al., 2023) to generate a non-backdoor perturbation (“evalCodeoOpenraught”) at the first position for the code snippet. Then, we inject this perturbation into different identifier positions of the code snippet and test the model’s predictions, plotting the results as the blue line in Figure 5. Likely, we inject the factual backdoor trigger (“testo_init”) into different identifier positions of the code snippet and test the model’s predictions, plotting the results as the red line in Figure 5. Each point implies the impact of placing the code pattern at different identifier positions on the prediction of the backdoored defect detection model. Both non-backdoor perturbations and backdoor triggers target the label “non-defective”, meaning that if “Probability” is less than 0.5, the attack is successful. This figure shows that only when the non-backdoor perturbation is injected at certain positions (e.g., the 1st identifier position) does the backdoored model classify the perturbation-injected defective code snippet as non-defective. In contrast, the backdoored model classifies the defective trigger-injected code snippet as non-defective, regardless of where the backdoor trigger is injected. It means that the robustness of the backdoor trigger is higher than that of the non-backdoor perturbation. In other words, the backdoored model is very sensitive to the backdoor trigger, regardless of its injection position. Therefore, intuitively, we can inject randomly initialized triggers at any identifier position for optimization. However, the backdoored model is not sensitive to non-backdoor perturbations, but injecting the randomly initialized trigger at certain positions is more likely to optimize effective non-backdoor perturbations rather than effective backdoor triggers. Therefore, if we can identify which positions are more likely to produce non-backdoor perturbations, we can inject the randomly initialized trigger into positions other than these to exclude the interference of non-backdoor perturbations, thereby improving trigger optimization efficiency.
@@
To this end, we investigate the sensitivity of the backdoored model to changes at each identifier position in the code. Specifically, we analyze the model’s sensitivity to each identifier position by masking each position in the code snippet and then calculating the loss value for predicting the masked code snippet as the ground-truth label. In Figure 5, the black dashed line represents the loss value of the backdoored model predicting the original code snippet as the ground-truth label. We also plot the loss values of the backdoored model predicting each masked code snippet as the ground-truth label as the orange line in Figure 5. The larger the change in loss value (the farther the orange triangle is from the black dashed line), the more sensitive the model is to the variation at that identifier position. From Figure 5, it is observed that sensitive identifier positions, such as the 1st, 2nd, and 8th identifier positions, are likely to produce effective non-backdoor perturbations. Compared to non-backdoor perturbations, the generation of the effective backdoor trigger is less correlated with the sensitivity of each position. Therefore, we can inject the randomly initialized trigger in insensitive identifier positions for optimization to reduce the probability of generating effective non-backdoor perturbations instead of effective backdoor triggers during the optimization process, thus improving trigger optimization efficiency. For instance, Figure 5 shows the distribution of the number of identifiers in code snippets of all clean samples. Observe that the number of identifiers in different code snippets varies, with most code snippets containing only a few identifiers. We experiment with optimizing the randomly initialized trigger injected into the top-ranked less sensitive positions covering the majority of code snippets. The backdoored defect detection model involved in the experiment is built on CodeBERT, and the experimental results are shown in Figure 5. Observe that injecting randomly initialized triggers at the least sensitive positions of each code snippet requires only 25 epochs to optimize an effective backdoor trigger, while more sensitive positions require more epochs. Some positions, such as the 4th least sensitive position from the end, do not even yield an effective trigger after 100 epochs of searching.
@@ 4.4. GCG-based Trigger Inversion
The Greedy Coordinate Gradient (GCG) proposed by Zou et al. (Zou et al., 2023) is used to search for an adversarial suffix onto the user prompt, which is intended to induce LLMs to respond to the user’s original, potentially harmful, request, i.e., producing undesirable behavior. Such a suffix can also be viewed as an adversarial perturbation. In our scenario, EliBadCode aims to search for a backdoor trigger/perturbation injected into the code snippet, which is intended to induce the backdoored NCM to produce target behaviors. Therefore, we borrow GCG to implement trigger inversion. As mentioned in Section 4.3, not every adversarial perturbation is an attacker-crafted backdoor trigger. Our goal is to eliminate backdoor triggers rather than non-backdoor perturbations. Therefore, unlike (Zou et al., 2023) which does not need to care about the suffix injection position, EliBadCode injects the trigger into positions where the NCM is less sensitive (i.e., the masked identifier positions in the selected masked samples), to reduce the interference of non-backdoor perturbations.
@@
Input: $X^{m}$, $Y$ selected masked samples, labels
$V$, $f_{\theta^{*}}$ trigger vocabulary, backdoored NCM
$\epsilon$, $k$, $r$ the times of iterations, the number of candidate substitutes, the times of repeat, respectively
$\beta$ the threshold for trigger anchoring
@@
3:    $\bm{e}_{S^{m}}\leftarrow$ produce embeddings of codes in $S^{m}$ using $f_{\theta^{*}}$
@@
6:       $\bm{e}_{t}\leftarrow$ produce $o_{t}$’s embeddings using $f_{\theta^{*}}$
@@
8:       $G\leftarrow\nabla o_{t}\mathcal{L}(f_{\theta^{*}}(\bm{e}^{\prime}_{S^{m}}),y^{\prime})$
9:       $\mathcal{T}\leftarrow$ select substitutes for each trigger token based on the top-$k$ gradients of $o_{t}$ in $G$
@@
14:          $\mathcal{T}_{i}\leftarrow$ get all substitutes for the $i$-th token of $t^{j}$
@@
21:    $l\leftarrow\mathcal{L}(f_{\theta^{*}}(S^{m}\oplus t),y^{\prime})$
@@
24:function TriggerAnchoring($S^{m}$,$t$, $y^{*}$)
25:    $t^{*}\leftarrow\emptyset$
26:    $l\leftarrow\mathcal{L}(f_{\theta^{*}}(S^{m}\oplus t),y^{*})$
27:    for each token $t_{i}$ in $t$ do
28:       $l_{i}\leftarrow\mathcal{L}(f_{\theta^{*}}(S^{m}\oplus(t\setminus t_{i})),y^{*})$
29:       if $|l-l_{i}|>\beta$ then
30:          $t^{*}\leftarrow t^{*}\cup t_{i}$
31:       end if
32:    end for
33:    return $t^{*}$
34:end function
35:
[…]
@@
44:$y^{*}$, $t\leftarrow$ run the outlier detection on $l^{C}$ and $t^{C}$ to detect the target label $y^{*}$ and the corresponding trigger $t$
@@
Algorithm 1 illustrates the GCG-based trigger inversion of EliBadCode in detail. In addition to the selected masked samples ($X^{m}$), trigger vocabulary ($V$), and a backdoored NCM ($f_{\theta^{*}}$) as shown in Figure 2, EliBadCode takes as input the labels ($Y$) and some key settings including times of iterations ($\epsilon$), the number of candidate substitutes ($k$), times of repeat ($r$), and the threshold for trigger anchoring ($\beta$). To eliminate backdoors in NCMs, EliBadCode first obtains the possible target label $y^{\prime}$ from $Y$ and gets masked code snippets $S^{m}$ with the label $y^{\prime}$ from $X^{m}$, then invokes the TriggerInversion function (lines 38–40). Then, in the TriggerInversion function, EliBadCode first randomly initializes a trigger ($t$) with $n$ tokens using $V$ (line 2), and then transforms code snippets in $S^{m}$ into vector representations (also called embeddings) $\bm{e}_{S^{m}}$ using the embedding layer of $f_{\theta^{*}}$ (line 3). Based on $\bm{e}_{S^{m}}$, it further iteratively optimizes $t$ $\epsilon$ times (lines 4–20). During each iteration, EliBadCode first generates the one-hot representation of $t$, denoted as $o_{t}$ (line 5). Second, it produces the embeddings of $o_{t}$ using $f_{\theta^{*}}$, denoted as $\bm{e}_{t}$ (line 6). Third, it injects $\bm{e}_{t}$ into $\bm{e}_{S^{m}}$ to produce the embeddings of trigger-injected masked code snippets, denoted as $\bm{e}^{\prime}_{S^{m}}$ (line 7). Forth, it feeds $\bm{e}^{\prime}_{S^{m}}$ to $f_{\theta^{*}}$ to compute gradients for $o_{t}$, denoted as $G$ (line 8). Fifth, based on the top-$k$ negative gradients of each trigger token in $G$, it selects substitutes for all trigger tokens in $t$, denoted as $\mathcal{T}$ (line 9). Based on $\mathcal{T}$, it generates a set of candidate triggers $T^{C}$ by repeating $r$ times, each time randomly replacing one token in $t$ with a random substitute in $\mathcal{T}$ (lines 10–18). Sixth, it injects each candidate trigger into $S^{m}$, calculates the loss values $l$ of $f_{\theta^{*}}$ predicting the trigger-injected code snippets as $y^{\prime}$, and selects the candidate trigger resulting in the smallest loss value as the inverted trigger (line 19). Finally, it calculates the loss value $l$ about the inverted trigger $t$ and the possible target label $y^{\prime}$ and returns them (lines 21–22). After iterating over all possible target labels and producing a set of loss values and the corresponding inverted triggers, one for each label. EliBadCode runs the outlier detection method (Wang et al., 2019) to obtain the ground-truth target label $y^{*}$ and the corresponding inverted trigger $t$. Next, $y^{*}$ and the corresponding $t$ will be input into the TriggerAnchoring function to obtain the effective components in $t$ (line 45).